A nonionizing solid dissolved in water changed the freezing point to -2.79°C. Calculate its concentration.


1
Expert's answer
2022-05-06T06:38:02-0400

2.79=1.86 K kg mol−1*m(solid), m(solid)=2.79/1.86=1.5 kg